Circular No. 6006

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GX 339–4 R. Sood, Australian Defence Force Academy, Canberra; and D. Camp-

bell-Wilson, Molonglo Radio Observatory, University of Sydney, communicate: “We report the detection of radio emission from the x-ray source GX 339–4 (4U 1658–48). Observations of this previously-radio-quiet black-hole candidate were made with the Molonglo Observatory Synthesis Telescope at 843 MHz, imaging a 23′ × 31′ field with 44′′ resolution. The timeaveraged flux between May 25.35 and 25.85 UT was 10.0 ± 0.7 rms mJy. The observations were repeated during the following week with these results: June 1.35–1.60, 6 ± 0.9 rms mJy; 2.41–2.60, 7.0 ± 1.3; 3.35–3.84, 6.8 ± 0.7. Preliminary data analysis does not reveal, within the constraints of our resolution, the extended radio jet structure for GX 339–4 predicted by Miyamoto and Kitamoto (1991, Ap.J. 374, 741).”

COMET TAKAMIZAWA-LEVY (1994f)

Extension to ephemeris on IAUC 5997:

1994 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 12 13^{h}30.^{m}93 +65^{o}19.^{′}9 1.117 1.392 81.^{o}4 46.^{o}1  8.7
17 12 46.26 +60 32.9 1.225 1.410 77.4 44.7  8.9
22 12 19.88 +55 58.8 1.344 1.431 73.2 42.8  9.2
27 12 03.43 +51 55.1 1.469 1.455 68.9 40.7  9.5
July 2 11 52.70 +48 23.0 1.597 1.483 64.7 38.3  9.7
7 11 45.52 +45 19.1 1.727 1.513 60.6 35.8 10.0
12 11 40.67 +42 39.1 1.854 1.546 56.5 33.2 10.2
17 11 37.42 +40 19.0 1.979 1.581 52.5 30.6 10.5
22 11 35.30 +38 15.6 2.100 1.618 48.5 28.1 10.7
27 11 34.03 +36 26.0 2.217 1.657 44.7 25.5 10.9
Aug. 1 11 33.38 +34 47.9 2.328 1.698 41.0 23.1 11.1
6 11 33.23 +33 19.7 2.433 1.740 37.5 20.8 11.3

Total visual magnitude estimates (cf. IAUC 5992): May 29.94 UT, 8.4 (A. Diepvens, Balen, Belgium, 0.15-m refractor); June 1.90, 8.7 (K. Sárneczky, Budapest, Hungary, 20×60 binoculars); 8.27, 8.0 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

1994 June 9 Daniel W. E. Green